F=ax+bt.t+c where F is force,x is distance and t is time. What is the dimension of axc/bt.t?

Grade:

1 Answers

Sanjeev Malik IIT BHU
12 years ago

try it to practise yourself. see the hint:

as F = ax + bt.t + c

this means that individually ax  , bt.t and c has unit of force i.e N.

ax= dimension of force

bt.t= "    "    "   "    " 

c=  "   "    "      "     "

clearly putting in que,

(in dimension form)  = F*F/F    = F  i.e ans is dimension of force itself.

 

 

please approve by clicking yes below